Find the force necessary to keep the toy car in equilibrium. The mass of the toy is 5kg and the inclination of the plane is 30°. Friction is neglected.

1
Expert's answer
2021-11-09T10:58:36-0500

Given:

"\theta=30^{\circ}"

"m=5\:\rm kg"


The force required to keep the toy car in equilibrium

"F=mg\sin\theta=5.0*9.8*\sin 30^{\circ}=25\:\rm N"